OpenStudy (meilendurcer):

1/3 divide 5/6= A. 2/5 B. 5/2 C. 2/15 D. 5/18

OpenStudy (anonymous):

start with \[\frac{1}{3}\times \frac{6}{5}\]

OpenStudy (meilendurcer):

i did

OpenStudy (anonymous):

did you get \[\frac{2}{5}\]?

OpenStudy (meilendurcer):

yes